# GAL-MAD: Towards Explainable Anomaly Detection in Microservice Applications Using Graph Attention Networks

### v1

@@ 3.2 GAL-MAD Model
Figure 5: GAT layer of the GAL-MAD
The decoder mirrors the encoder’s structure, comprising a one-to-many LSTM followed by two GAT layers, to reconstruct along the time axis and feature axis respectively. Since GAT cannot output the required three-dimensional shape with the exact number of output features in the reconstruction, due to the incompatibility of the number of outputs and the number of attention heads, we employed a linear layer at all output nodes to adjust the feature axis to the desired size.
Normal data was used to train the model, which led to the expected higher MSE for anomalous data during testing. At the end of the reconstruction, a sigmoid function was used to get the final prediction. The model was trained using a batch size of 360 and a learning rate of 0.001 with a decay of 0.5 per epoch using the Adam optimizer.
@@ 3.3 Anomalous Service Localization
Anomalous service localization was conducted using SHAP on anomalous model detections. The results are depicted in Table 4. The process of locating the anomalous service is as follows,
  1. 1.
    SHAP values are calculated for each time step of a period
  2. 2.
    Anomalously detected time steps are filtered from the period
  3. 3.
    SHAP values are aggregated along the time axis
  4. 4.
    Absolute sum is computed over the resulting matrix along the feature axis
    1. (a)
      The SHAP values for the response time feature and container_fs_usage_bytes are weighted 4 times to improve results
  5. 5.
    Service is deduced by the index of the maximum value in the resulting array
  6. 6.
    The anomalous feature is deduced by the index of the highest value in the respective service array
4 Results
@@ 4.1 Experiments on RS-Anomic Dataset
We evaluated the RS-Anomic dataset using state-of-the-art anomaly detection models, GDN Deng and Hooi (2021), MAD-GAN Li et al. (2019), Kitsune Mirsky et al. (2018) and a custom transformer-based reconstruction model for reference Vaswani et al. (2017). GDN uses cosine similarity to create a graph structure, predicts future values using graph attention-based forecasting, and calculates the graph deviation score. MAD-GAN is designed to detect anomalies in time series data and uses an LSTM-RNN discriminator and a reconstruction-based method. Kitsune is an unsupervised Network Intrusion Detection System that can detect local network attacks without supervision. We categorized anomalies as the positive class and compared the models using Accuracy (A), Recall (R), and Specificity (S) metrics. The normal data was split into an 80:20 ratio for training and testing. Only normal data was used to train models based on respective papers. We created three testing scenarios with normal:anomalous behaviour ratios of 95:5, 90:10, and 60:40. For each testing scenario, the test data was mixed with anomalous data, according to the respective ratios. Anomalous samples were balanced across different anomaly types. The difference in the number of response time features for each microservice, due to varying numbers of links between microservices, was unified by summing all recorded response times for each service to create one response time feature. If a service experienced a network delay in any communication link, this information was preserved by summing the response times. In addition to response times, we introduced a moving average of response times for 5 minute and 30 minute windows. The input dimensions for all the models were 264 for each time step assembled by concatenating 22 features, including moving averages for 12 services. Moreover, standard scaling to scale our data based on normal data distribution. We considered a window size of 24 data points. We trained all models for 20 epochs with Adam optimizer and learning rates 0.001, 0.0005, 0.1 and 0.05 for GDN, MAD-GANs, Kitsune and Transformer, respectively. The results for GDN, MAD-GAN, Vanilla transformer, Kitsune and GAL-MAD models on the RS-Anomic dataset are displayed in Table 2. According to the results, especially the recall value, state-of-the-art anomaly detection models are not very successful in detecting anomalies in microservices. This indicates that RS-Anomic is a challenging dataset, and further research is necessary to develop new Machine Learning models for microservice anomaly detection.
@@
The RS-Anomic dataset includes performance metrics of a microservice application under both normal and anomalous behaviour. This dataset enables anomaly detection in microservice applications and allows for the development of an anomaly classification model, as it contains ten different types of anomalies. However, it is essential to note that our application was deployed on a single server, therefore, the RS-Anomic dataset does not capture the impacts of a distributed deployment. Most microservice applications operate in distributed environments with auto-scaling, which introduces additional complexities that are not reflected in the RS-Anomic dataset. Table 2 presents the results of our empirical study, evaluating the existing anomaly detection models using the newly introduced RS-Anomic dataset demonstrates that there is still room to develop more accurate anomaly detection models. The suboptimal performance of current models may be attributed to their inability to account for the graph-like structure inherent in the RS-Anomic dataset. This structure, derived from the architecture of microservice applications, is critical for identifying dependencies among input features.
Model 95:5 90:10 60:40
A R S A R S A R S
GDN 0.9961 0.8093 0.9645 0.9409 0.7684 0.9620 0.8361 0.7766 0.8741
MAD-GANs 0.6073 0.8627 0.5920 0.5669 0.6893 0.5521 0.5225 0.5238 0.5258
Kitsune 0.9757 0.9006 0.9804 0.8984 0.5529 0.9412 0.5921 0.4142 0.7077
Transformer 0.9810 0.8252 0.9905 0.9678 0.7994 0.9885 0.8712 0.7491 0.9506
GAL-MAD 0.9838 0.9884 0.9559 0.9788 0.9228 0.9855 0.9500 0.8938 0.9859
@@ A R S A R S A R S
Table 2: Performance of GAL-MAD and the state-of-the-art anomaly detection models on RS-Anomic dataset in terms of Accuracy(A), Recall(R), and Specificity(S) testing under a mixture of normal to anomaly ratios of 95:5, 90:10 and 60:40
4.2 Experiments on GAL-MAD
The GAL-MAD detector predicts anomalies in 80-second windows. Normal data windows produce a loss lower than 2.0. The reconstruction loss of the autoencoder model is passed through a sigmoid function to obtain the prediction as follows:
$prediction=sigmoid(loss-c)$
Where $loss$ is the reconstruction loss and $c$ is the upper bound of the reconstruction loss observed for normal data. Table 3 shows average loss values for two configurations of the model, one using only cAdvisor metrics and the other using both cAdvisor metrics and response time features. The response time features include moving averages calculated over 5 minute and 30 minute windows. The analysis reveals the significant impact of incorporating response time features on anomaly detection, particularly for network-related anomalies. Moreover, packet loss, out-of-order packets, and low bandwidth anomalies were tested under two different strengths to highlight the lack of sensitivity observed on fine network anomalies.
@@ 4.2 Experiments on GAL-MAD
Table 3: Losses with and without response time for each anomaly type
4.2.1 Comparative Study of GAL-MAD Model
We conducted a comparative study with the state of the art anomaly detection models on our dataset. The results in Table 2 show that our model has better performance, given the complex nature of the data.
@@ 4.3 Anomalous Service Localization Scores
Anomaly localizations are considered true when both the service with the injected anomaly and its related features are successfully identified. SHAP values visualized as heatmaps for selected anomaly instances are shown in Figures 6 to 9. The features are indicated by labels 0 to 21 in the X axis, listed in order as shown in Table 1, followed by the moving averages of response times in 5 minute and 30 minute windows labeled as 20 and 21, respectively. The Y axis represents the payment, shipping, redis, mongodb, dispatch, rabbitmq, user, mysql, catalog, ratings, web and cart services in order.
The heatmaps provide intuitive explanations in cases related to performance and response time anomalies. Figure 6 illustrates a heatmap with diverging shap values for response time features when the API calls from the catalogue service to the mongodb service were delayed. Figure 7 demonstrates that CPU usage related features of the dispatch service with a diverging SHAP value when an abnormally high CPU usage is introduced.
However, it is notable that fine network-related anomalies such as out-of-order-packets and packet loss are substantially more difficult to deduce from visualizations Figures 8 and 9. The SHAP values were modified by increasing the weight of container_fs and response times features fourfold, with results documented in Table 4.
@@ 5 Conclusion
This research introduces the RS-Anomic dataset, a comprehensive multivariate time-series resource designed to advance anomaly detection research in microservices architectures. The dataset encompasses 100,000 normal data points and 14,000 anomalous data points, spanning ten distinct anomaly types, providing a robust foundation for academic and industrial researchers. Furthermore, we propose the GAL-MAD model for anomaly detection, which leverages Graph Attention Networks and Long Short-Term Memory networks to capture both spatial and temporal dependencies between microservices. Our empirical study evaluated existing anomaly detection models using the RS-Anomic dataset. The GAL-MAD model outperformed the state-of-the-art anomaly detection on the RS-Anomic dataset. The results of using SHAP for explainability show that it facilitates localizing the anomalous service and the feature to a great extent in performance-related anomalies. System administrators can leverage such a framework to quickly identify and diagnose anomalies, understand the root cause of performance degradations and, gain deeper insight to microservice interactions. The RS-Anomic dataset and GAL-MAD model represent important steps toward more intelligent and interpretable anomaly detection in complex distributed systems. Further research directions include expanding the dataset to incorporate multi-node architectures and auto-scaling capabilities that more closely reflect production-level microservices deployments. Additionally, while SHAP provides insights into root cause localization, further work is required in interpreting anomalies by integrating the application architecture context. Further analysis of network-related metrics is required to improve network anomaly explainability and diagnostic precision.
Figure 6: Heat map for response time delay anomaly in the mongo database service. Service 3 represents the MongoDB service. Features 19, 20, and 21 represent response times and moving averages of the response times for 5 and 30 minute windows, respectively.
Figure 7: Heat map for high CPU usage in the dispatch service anomaly. Service 4 represents the dispatch service. Features 1,2,3, and 5 represent memory usage, memory failures, memory working set usage and CPU usage time, respectively.
Figure 8: Heat map for a false localization of out-of-order packets in the user service.
Figure 9: Heat map for a false localization of packet loss anomaly in the user service.

### v2

@@ 3.2 Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD) Model
The GAL-MAD decoder reconstructs the input by mirroring the encoder’s architecture, which comprises a one-to-many LSTM followed by two GAT layers.
The GAL-MAD model is trained on normal data, using MSE. Let the reconstruction error of GAL-MAD be $\mathcal{L}$. During inference, a higher MSE means the input is anomalous. The Sigmoid function is applied to MSE during inference to derive the final prediction as follows:
$y=Sigmoid(\mathcal{L}-c)$ (7)
, where $c$ is the upper bound of the reconstruction loss observed for normal data. In our experiments, we set $c=2.0$.
3.2.1 Implementation Details
The model was trained using the Adam optimiser with a batch size of 360, a learning rate of 0.001, and a decay of 0.5 per epoch. Each LSTM uses a sequence length 24, translating to 2-minute time windows polled every 5 seconds. Further, we set $d_{z}$ to 1. The GAL-MAD predicts anomalies in 80-second windows.
@@ 3.3 Anomalous Service Localization
To enhance the interpretability of anomaly detection in microservice architectures, we employ SHapley Additive exPlanations (SHAP) Lundberg and Lee (2017a) to localize anomalous services and identify contributing features. Upon detecting an anomaly, SHAP values are computed for each time step within the observation period, quantifying the contribution of each feature to the model’s output. Time steps identified as anomalous by the detection model are selected for further analysis. The SHAP values corresponding to these anomalous time steps are aggregated along the temporal axis to obtain a consolidated view of feature contributions. To emphasize critical features, specific SHAP values are weighted; in particular, the SHAP values for the $responsetime$ and $container\_fs\_usage\_bytes$ features are multiplied by a factor of four. Subsequently, the absolute SHAP values are summed across features for each service to determine the overall contribution, with the service corresponding to the maximum sum identified as the most likely source of the anomaly. Within this anomalous service, the feature with the highest absolute SHAP value is determined, pinpointing the specific metric most responsible for the detected anomaly. This methodology facilitates the pinpointing of both the anomalous service and the specific feature contributing to the detected anomaly, thereby enhancing the explainability and actionable insights derived from the anomaly detection model.
4 Experimental Study
@@ 4.1 Experiments on RS-Anomic Dataset
We evaluated the RS-Anomic dataset using state-of-the-art anomaly detection models, GDN Deng and Hooi (2021), MAD-GAN Li et al. (2019), Kitsune Mirsky et al. (2018) and a Transformer-based reconstruction model Vaswani et al. (2017). GDN uses cosine similarity to create a graph structure, predicts future values using graph attention-based forecasting, and calculates the graph deviation score. MAD-GAN is designed to detect anomalies in time series data and uses an LSTM-RNN discriminator and a reconstruction based method. Kitsune is an unsupervised Network Intrusion Detection System that can detect local network attacks without supervision. We categorized anomalies as the positive class and compared the models using Accuracy (A), Recall (R), and Specificity (S) metrics. The normal data was split into an 80:20 ratio for training and testing. Only normal data was used to train models based on respective papers. We created three testing scenarios with normal:anomalous behaviour ratios of 95:5, 90:10, and 60:40. For each testing scenario, the test data was mixed with anomalous data, according to the respective ratios. Anomalous samples were balanced across different anomaly types. The difference in the number of response time features for each microservice, due to varying numbers of links between microservices, was unified by summing all recorded response times for each service to create one response time feature. If a service experienced a network delay in any communication link, this information was preserved by summing the response times. In addition to response times, we introduced a moving average of response times for 5 minute and 30 minute windows. The input dimensions for all the models were 264 for each time step assembled by concatenating 22 features, including moving averages for 12 services. Moreover, standard scaling to scale our data based on normal data distribution. We considered a window size of 24 data points. We trained all models for 20 epochs with Adam optimizer and learning rates 0.001, 0.0005, 0.1 and 0.05 for GDN, MAD-GANs, Kitsune and Transformer, respectively. The results for GDN, MAD-GAN, Vanilla transformer, Kitsune and GAL-MAD models on the RS-Anomic dataset are displayed in Table 2. According to the results, especially the recall value, state-of-the-art anomaly detection models are not very successful in detecting anomalies in microservices. This indicates that RS-Anomic is a challenging dataset, and further research is necessary to develop new Machine Learning models for microservice anomaly detection. The suboptimal performance of current models may be attributed to their inability to account for the graph-like structure inherent in the RS-Anomic dataset. This structure, derived from the architecture of microservice applications, is critical for identifying dependencies among input features.
@@
While the RS-Anomic dataset enables anomaly detection in microservice applications and allows for the development of an anomaly classification model, it is essential to note that our application was deployed on a single server. Therefore, the RS-Anomic dataset does not capture the impacts of a distributed deployment. Most microservice applications operate in distributed environments with auto-scaling, which introduces additional complexities not reflected in the RS-Anomic dataset.
Ratio Metric GDN MAD-GAN Kitsune Transformer GAL-MAD
95:5 A 0.9961 0.6073 0.9757 0.9810 0.9838
R 0.8093 0.8627 0.9006 0.8252 0.9884
S 0.9645 0.5920 0.9804 0.9905 0.9559
90:10 A 0.9409 0.5669 0.8984 0.9678 0.9788
R 0.7684 0.6893 0.5529 0.7994 0.9228
S 0.9620 0.5521 0.9412 0.9885 0.9855
60:40 A 0.8361 0.5225 0.5921 0.8712 0.9500
R 0.7766 0.5238 0.4142 0.7491 0.8938
S 0.8741 0.5258 0.7077 0.9506 0.9859
@@
Table 2: Performance of GAL-MAD and state-of-the-art anomaly detection models on the RS-Anomic dataset across varying anomaly-to-normal data ratios (95:5, 90:10, 60:40), measured in terms of Accuracy (A), Recall (R), and Specificity (S).
@@
Table 3: Reconstruction loss with and without response time for each anomaly type
4.2 Experiments on GAL-MAD
We first conduct a comparative study with the state-of-the-art anomaly detection models on the newly introduced RS-Anomic dataset. The results in Table 2 show that our model has achieved better performance across all the ratios. In particular, for all the ratios, the proposed GAL-MAD has achieved the highest recall compared to the existing anomaly detection methods, indicating the effectiveness of GAL-MAD in detecting anomalies. In anomaly detection, particularly within imbalanced datasets where anomalies are rare, recall (also known as sensitivity) is often the most critical metric. Recall measures the proportion of actual anomalies correctly identified by the model, ensuring that true anomalies are not overlooked. Missing an anomaly (a false negative) can have significant consequences.
Next, we evaluate the effect of including the response time as a feature in detecting the anomalies. Table 3 shows average loss values for the two configurations, one using only cAdvisor metrics and the other using both cAdvisor metrics and response time features. The response time features include moving averages calculated over 5-minute and 30-minute windows. The analysis reveals the significant impact of incorporating response time features on anomaly detection, particularly for network-related anomalies. Moreover, packet loss, out-of-order packets, and low bandwidth anomalies were tested under two different strengths to highlight the lack of sensitivity observed on fine network anomalies.
Finally, we conduct an ablation study to evaluate the impact of each component of the proposed GAL-MAD. Table 4 shows the results of the ablation study. The LSTM component was removed from the encoder and the decoder of the GAL-MAD detector to create the GAT-Autoencoder (GAT-AE), and both GAT layers were removed from the encoder and the decoder of the GAL-MAD detector to make the LSTM-Autoencoder (LSTM-AE). Finally, the GAT and the LSTM components were removed to construct a linear autoencoder (Linear-AE). The results demonstrate that the proposed GAL-MAD achieves the highest recall for all scenarios, indicating the importance of each component of the proposed architecture.
Ratio Metric GAL-MAD GAT-AE LSTM-AE Linear-AE
95:5 A 0.9838 0.9808 0.9897 0.9893
R 0.9884 0.8839 0.9032 0.8903
S 0.9559 0.9867 0.9947 0.9953
90:10 A 0.9788 0.9704 0.9784 0.9784
R 0.9228 0.8460 0.8492 0.8892
S 0.9855 0.9867 0.9953 0.9953
60:40 A 0.9500 0.9264 0.9371 0.9358
R 0.8938 0.8333 0.8462 0.8425
S 0.9859 0.9857 0.9953 0.9955
Table 4: Ablation study of GAL-MAD Model in terms of Accuracy (A), Recall (R), and Specificity (S) for normal to anomaly ratios of 95:5, 90:10, and 60:40. AE stands for Auto Encoder.
@@ 4.3 Anomalous Service Localization Scores
(a) Response time delay anomaly in the MongoDB service
(b) High CPU usage in the Dispatch service anomaly
(c) Out-of-order packets in the User service anomaly
(d) Packet loss anomaly in the User service
Figure 4: Heatmaps illustrating various anomalies detected in microservices
As described in Section 3.3, we use eXplainable Artificial Intelligence (XAI) technique, SHAP Lundberg and Lee (2017a), to localize the anomalous service once an anomaly is detected. SHAP values were computed and visualized as heatmaps for selected anomalous instances, as depicted in Figures 4(a) to 4(d). In these heatmaps, the x-axis represents features (performance metrics) labelled from 0 to 21, corresponding to the sequence outlined in Table 1. Specifically, labels 20 and 21 denote the moving averages of response times over 5-minute and 30-minute intervals, respectively. The y-axis enumerates the microservices in the following order: payment, shipping, redis, mongodb, dispatch, rabbitmq, user, mysql, catalog, ratings, web, and cart.
The heatmaps offer intuitive insights, particularly for anomalies related to performance and response times. For instance, Figure 4(a) illustrates divergent SHAP values for response time features when API calls from the catalog service to the mongodb service experienced delays. Similarly, Figure 4(b) highlights significant SHAP values associated with CPU usage features in the dispatch service during instances of artificially induced high CPU utilization. Conversely, non-critical network-related anomalies, such as out-of-order packets and packet loss, pose challenges for visual interpretation. This is evident in Figures 4(c) and 4(d), where the SHAP value distributions are less distinct.
To enhance the detection of such anomalies, the SHAP values for the $container\_fs$ and response time features were amplified by a factor of four. The outcomes of this adjustment are documented in Table 5. In this table, an anomaly localization is deemed accurate when both the microservice experiencing the injected anomaly and its associated features are correctly identified.
@@ 5 Conclusion
This study introduces the RS-Anomic dataset, a comprehensive multivariate time-series resource designed to advance anomaly detection research in microservices architectures. The dataset comprises 100,000 normal and 14,000 anomalous data points, encompassing ten distinct anomaly types, providing a robust foundation for academic and industrial research endeavors. Building upon this dataset, we propose the Graph Attention and LSTM-based Microservice Anomaly Detection (GAL-MAD) model, which leverages Graph Attention Networks and Long Short-Term Memory networks to capture spatial and temporal dependencies among microservices. Empirical evaluations demonstrate that GAL-MAD outperforms existing state-of-the-art anomaly detection models on the RS-Anomic dataset, achieving superior recall across varying anomaly ratios. To enhance interpretability, SHAP values were employed, facilitating the localization of anomalous services and identifying contributing features, particularly in performance-related anomalies. This explainability empowers system administrators to diagnose anomalies swiftly, comprehend the root causes of performance degradations, and gain deeper insights into microservice interactions.
The RS-Anomic dataset and the GAL-MAD model represent significant strides toward intelligent and interpretable anomaly detection in complex distributed systems. Future research directions include expanding the dataset to incorporate multi-node architectures and auto-scaling capabilities that more accurately reflect production-level microservices deployments. Additionally, while SHAP provides valuable insights into root cause localization, further work is warranted to interpret anomalies within the context of application architecture. Moreover, in-depth analysis of network-related metrics is essential to enhance the explainability and diagnostic precision of network anomalies.